# Non-oscillatory entropy stable DG schemes for hyperbolic conservation law

### v1

@@
Abstract: In this paper, we propose a class of non-oscillatory, entropy-stable discontinuous Galerkin (NOES-DG) schemes for solving hyperbolic conservation laws. By incorporating a specific form of artificial viscosity, our new scheme directly controls entropy production and suppresses spurious oscillations. To address the stiffness introduced by the artificial terms, which can restrict severely time step sizes, we employ the integration factor strong stability-preserving Runge-Kutta method for time discretization. Furthermore, our method remains compatible with positivity-preserving limiters under suitable CFL conditions in extreme cases. Various numerical examples demonstrate the efficiency of the proposed scheme, showing that it maintains high-order accuracy in smooth regions and avoids spurious oscillations near discontinuities.
@@ 1 Introduction
In this work, motivated by [15], we develop a class of provably (in the semi-discrete sense) high order and entropy stable DG methods that preserve all desired properties of the classic DG method and is capable of eliminating spurious oscillations in the presence of shocks. The difficulty lies in the design of an artificial term that can simultaneously control entropy production and oscillations, as well as preserve the original high order accuracy. Furthermore, it is also desired that the modified formulation is convenient to implement and cost-effective. To address this, we generalize the entropy-based artificial diffusion framework [18] by incorporating an idea from the essentially oscillation-free DG (OFDG) method [32]. The OFDG method employs a projection-based damping term in the DG formulation to control spurious oscillations. In particular, the damping coefficient is adaptively adjusted by measuring discontinuity intensity, decreasing in smooth regions and increasing near discontinuities. In this work, we employ an artificial diffusion term rather than a damping term, and furthermore, the coefficient is determined to ensure both the entropy stability and oscillation-free property. Our work introduces three key improvements over existing approaches. First, in [15], the authors claimed that an arbitrary piecewise polynomial approximation or interpolation with desired high order accuracy for the entropy variable can ensure the entropy dissipation property. However, we disprove this claim by providing a counterexample. Moreover, we demonstrate that a specific interpolation property is required for ensuring provable entropy stability. Second, as mentioned in [8], the denominator of the coefficient in the entropy correction term can approach zero more rapidly than the numerator, potentially compromising accuracy. To circumvent the difficulty, we propose to further incorporate the oscillation coefficient to control the main coefficient, thereby preserving the accuracy of the scheme. Last, the proposed method attempts to directly control entropy production, allowing it to accommodate arbitrary monotone numerical fluxes and quadrature rules. Moreover, the algorithm can be implemented by reusing a classic DG code with only the artificial viscosity term added, significantly reducing development time.
@@
Note that the proposed entropy artificial viscosity term may incur additional stiffness, which can lead to stringent CFL time step restriction when a standard explicit strong-stability-preserving Runge-Kutta (SSP-RK) time discretization is used. To address this issue, we implement the explicit integral factor SSP-RK time discretization [6], which can greatly mitigate the time step constraint. Furthermore, for extreme problems such as the high Mach number astrophysical jet problem, it is necessary to ensure positivity of density and pressure to avoid simulation failures. Note that, as the artificial viscosity term is locally conservative, we can further adapt the positivity-preserving limiter developed in [41] with an appropriate adjustment on the CFL condition.
@@
The rest of this paper is organized as follows. In Section 2, we introduce the formulation of semi-discrete non-oscillatory entropy stable DG schemes for both one-dimensional and two-dimensional scalar equations and systems, along with theoretical analysis on mass conservation, discrete entropy stability, and error estimation. In section 3, we present the fully discrete formulations, integrating an integral factor SSP-RK time discretization and a modified positivity-preserving limiter. In Section 4, various numerical examples are provided to demonstrate the efficiency and efficacy of our proposed scheme. The conclusions are discussed in Section 5.
@@ 2.1.1 The classical DG discretization
where $P^{k}(K)$ is the space of polynomials of degree at most $k$ on cell $K$. Notice that for any $w\in V_{h}^{k}$, it can be discontinuous at the cell boundary $x_{i\pm 1/2}$. In the following, we let $w^{+}_{i+1/2}$ (resp. $w^{-}_{i+1/2}$) denote the limiting value of $w$ at $x_{i+1/2}$ from the element $K_{i+1}$ (resp. $K_{i}$), $[w]_{i+1/2}=w^{+}_{i+1/2}-w^{-}_{i+1/2}$ denote its jump, and $\{w\}_{i+1/2}=\frac{1}{2}(w^{+}_{i+1/2}+w^{-}_{i+1/2})$ be its average at $x_{i+1/2}$.
@@ 2.1.2 The entropy stable non-oscillatory DG schemes
$\sigma_{i}^{jump}$ $=c_{f}\left(h\left\|\left[u_{h}\right]\right\|_{\partial K_{i}}+\sum_{l=1}^{k}{l\left(l+1\right)h^{l+1}\left\|\left[\partial_{x}^{l}u_{h}\right]\right\|_{\partial K_{i}}}\right),$ (2.12a)
@@
Meanwhile, the DG formulation (2.9) with $\sigma_{i}=\sigma^{entropy}_{i}$, while entropy stable, may still generate spurious oscillations in the presence of strong shocks. To remedy such a drawback, we further incorporate $\sigma_{i}^{jump}$ to suppress oscillations and let $\sigma_{i}=\max\left\{\sigma_{i}^{jump},\sigma_{i}^{entropy}\right\}$. The $\sigma_{i}^{jump}$, originally proposed by the OFDG method [32] and slightly modified in [39], measures discontinuity intensity and is defined with
$\left\|[\partial_{x}^{l}u_{h}]\right\|_{\partial K_{i}}=\left|\left[\frac{\partial^{l}u_{h}}{\partial x^{l}}\right]_{x_{i-1/2}}\right|+\left|\left[\frac{\partial^{l}u_{h}}{\partial x^{l}}\right]_{x_{i+1/2}}\right|.$
Note that $\sigma_{i}^{jump}$ is small in smooth regions, and becomes large near discontinuities.
@@
Here, $A\lesssim B$ means that there exists a constant $c_{0}>0$ independent of $h$ such that $A\leq c_{0}B$. However, we cannot directly estimate the ratio $F_{i}/E_{i}$ unless there is a strictly positive lower bound of $E_{i}$, which is not the case in practice. Therefore, in the simulation, we set
@@ 2.1.3 Theoretical properties
$\sigma_{i}^{entropy}<C\,\sigma_{i}^{jump},\quad\forall K_{i}\in\mathcal{K},$ (2.17)
@@
$\mathrm{RHS}$ $=-\sum_{i}{\left(\hat{f}_{i+1/2}^{D}v_{h}|_{i+1/2}^{-}-\hat{f}_{i+1/2}^{D}v_{h}|_{i+1/2}^{+}\right)}$ (2.20)
@@
$=-\sum_{i}{\frac{1}{2}\alpha_{i+1/2}\left(u_{h}|_{i+1/2}^{+}-u_{h}|_{i+1/2}^{-}\right)\left(v\left(u_{h}|_{i+1/2}^{+}\right)-v\left(u_{h}|_{i+1/2}^{-}\right)\right)}$
@@
$=$ $\sum_{i}\left(\sum_{\alpha=1}^{k+1}\omega_{\alpha}\frac{\partial u_{h}}{\partial t}(\hat{x}^{\alpha}_{i})v_{h}(\hat{x}^{\alpha}_{i})h+\mathcal{O}(h^{2k+1})\right)$
$=$ $\sum_{i}\left(\sum_{\alpha=1}^{k+1}\omega_{\alpha}\frac{\partial U(u_{h})}{\partial t}(\hat{x}^{\alpha}_{i})h+\mathcal{O}(h^{2k+1})\right)$
@@
Here, $\omega_{\alpha}$ and $\hat{x}^{\alpha}_{i}$ are the Gauss-Lobatto weights and quadrature points in each cell $K_{i}$, with $\hat{x}_{i}^{1}=x_{i-1/2}$ and $\hat{x}^{k+1}_{i}=x_{i+1/2}$. This indicates that the entropy can be considered stable, with only a small, high order growth. Also see [15].
@@
For $\sigma_{i}$, since $u$ is smooth, we have $\left|[\partial_{x}^{l}u_{h}]\right|=\left|[\partial_{x}^{l}(u-u_{h})]\right|$. According to (2.15), the viscosity coefficient has an upper bound $\sigma_{i}\leq C\sigma_{i}^{jump}$. Hence,
$\sum_{i}{\left(\sigma_{i}\right)^{2}}h^{-1}$ $\lesssim h\sum_{i}{\sum_{l=0}^{k}{\left\|\left[\partial_{x}^{l}u_{h}\right]\right\|_{\partial K_{i}}^{2}h^{2l}}}$
$=h\sum_{i}{\sum_{l=0}^{k}{\left\|\left[\partial_{x}^{l}\left(u-u_{h}\right)\right]\right\|_{\partial K_{i}}^{2}h^{2l}}}$
@@
$\sigma_{i,m}^{jump}$ $=c_{f}\left(h\left\|\left[u_{h,m}\right]\right\|_{\partial K_{i}}+\sum_{l=1}^{k}{l\left(l+1\right)h^{l+1}\left\|\left[\partial_{x}^{l}u_{h,m}\right]\right\|_{\partial K_{i}}}\right),$
@@
Here, we define $\sigma_{i}^{jump}$ measuring discontinuity on each component without characteristic decomposition. The numerical entropy flux $\hat{F}$ is given as the central flux (2.13) as well. The numerical flux $\hat{\mathbf{f}}=\hat{\mathbf{f}}(\mathbf{u}_{L},\mathbf{u}_{R})$ is splitted into the central part $\hat{\mathbf{f}}^{C}$ and the diffusive part $\hat{\mathbf{f}}^{D}$. For example, for the Lax-Friedrichs flux,
@@
Following the same idea in Theorem 2.2, we can also prove that scheme (2.31) is entropy stable for hyperbolic systems (2.28) with periodic or compact boundaries under the condition
$\sigma_{i}^{entropy}<C\max\limits_{m}\{\sigma_{i,m}^{jump}\},\quad\forall K_{i}\in\mathcal{K}.$
@@ 2.3 Two-dimensional system
$\sigma_{K,m}^{jump}$ $=c_{f}\left(h_{K}\left\|\left[u_{h,m}\right]\right\|_{\partial K}+\sum_{\left|l\right|=1}^{k}{\left|l\right|\left(\left|l\right|+1\right)h_{K}^{\left|l\right|+1}\left\|\left[\partial^{l}u_{h,m}\right]\right\|_{\partial K}}\right),\ l\in\mathbb{N}^{d},$

### v2

@@
Abstract: In this paper, we propose a class of non-oscillatory, entropy-stable discontinuous Galerkin (NOES-DG) schemes for solving hyperbolic conservation laws. By incorporating a specific form of artificial viscosity, our new scheme directly controls entropy production and suppresses spurious oscillations. To address the stiffness introduced by the artificial terms, which can restrict severely time step sizes, we employ the integrating factor strong stability-preserving Runge-Kutta method for time discretization. Furthermore, our method remains compatible with positivity-preserving limiters under suitable CFL conditions in extreme cases. Various numerical examples demonstrate the efficiency of the proposed scheme, showing that it maintains high-order accuracy in smooth regions and avoids spurious oscillations near discontinuities.
@@ 1 Introduction
In this work, motivated by [15], we develop a class of provably (in the semi-discrete sense) high order and entropy stable DG methods that preserve all desired properties of the classic DG method and is capable of eliminating spurious oscillations in the presence of shocks. The difficulty lies in the design of an artificial term that can simultaneously control entropy production and oscillations, as well as preserve the original high order accuracy. Furthermore, it is also desired that the modified formulation is convenient to implement and cost-effective. To address this, we generalize the entropy-based artificial diffusion framework [18] by incorporating an idea from the essentially oscillation-free DG (OFDG) method [32]. The OFDG method employs a projection-based damping term in the DG formulation to control spurious oscillations, and was extended in [33] to ensure a scale-invariant property. In addition, [33] showed that the damping technique, which damps the modal coefficients of the high-order moments of the numerical solution, is closely related to the artificial viscosity approach. In particular, the damping coefficient is adaptively adjusted by measuring discontinuity intensity, decreasing in smooth regions and increasing near discontinuities. In this work, we employ an artificial diffusion term rather than a damping term, and furthermore, the coefficient is determined to ensure both the entropy stability and oscillation-free property. Our work introduces three key improvements over existing approaches. First, in [15], the authors claimed that an arbitrary piecewise polynomial approximation or interpolation with desired high order accuracy for the entropy variable can ensure the entropy dissipation property. However, we disprove this claim by providing a counterexample. Moreover, we demonstrate that a specific interpolation property is required for ensuring provable entropy stability. Second, as mentioned in [8], the denominator of the coefficient in the entropy correction term can approach zero more rapidly than the numerator, potentially compromising accuracy. To circumvent the difficulty, we propose to further incorporate the oscillation coefficient to control the main coefficient, thereby preserving the accuracy of the scheme. Last, the proposed method attempts to directly control entropy production, allowing it to accommodate arbitrary monotone numerical fluxes and quadrature rules. Moreover, the algorithm can be implemented by reusing a classic DG code with only the artificial viscosity term added, significantly reducing development time.
@@
Note that the proposed entropy artificial viscosity term may incur additional stiffness, which can lead to stringent CFL time step restriction when a standard explicit strong-stability-preserving Runge-Kutta (SSP-RK) time discretization is used. To address this issue, we implement the explicit integral factor SSP-RK time discretization [6], which can greatly mitigate the time step constraint. Furthermore, for extreme problems such as the high Mach number astrophysical jet problem, it is necessary to ensure positivity of density and pressure to avoid simulation failures. Note that, as the artificial viscosity term is locally conservative, we can further adapt the positivity-preserving (PP) limiter developed in [42] with an appropriate adjustment on the CFL condition.
@@
The rest of this paper is organized as follows. In Section 2, we introduce the formulation of semi-discrete non-oscillatory entropy stable DG schemes for both one-dimensional and two-dimensional scalar equations and systems, along with theoretical analysis on mass conservation, discrete entropy stability, and error estimation. In Section 3, we present the fully discrete formulations, integrating an integral factor SSP-RK time discretization and a modified PP limiter. In Section 4, various numerical examples are provided to demonstrate the efficiency and efficacy of our proposed scheme. The conclusions are discussed in Section 5.
@@ 2.1.1 The classical DG discretization
where $P^{k}(K)$ is the space of polynomials of degree at most $k$ on cell $K$. Notice that for any $w\in V_{h}^{k}$, it can be discontinuous at the cell boundary $x_{i\pm 1/2}$. In the following, we let $w^{+}_{i+1/2}$ (resp. $w^{-}_{i+1/2}$) denote the limiting value of $w$ at $x_{i+1/2}$ from the element $K_{i+1}$ (resp. $K_{i}$), $[\![w]\!]_{i+1/2}=w^{+}_{i+1/2}-w^{-}_{i+1/2}$ denote its jump, and $\{w\}_{i+1/2}=\frac{1}{2}(w^{+}_{i+1/2}+w^{-}_{i+1/2})$ be its average at $x_{i+1/2}$.
@@ 2.1.2 The entropy stable non-oscillatory DG schemes
$\sigma_{i}^{jump}$ $=c_{f}\left(h\left\|[\![u_{h}]\!]\right\|_{\partial K_{i}}+\sum_{l=1}^{k}{l\left(l+1\right)h^{l+1}\left\|[\![\partial_{x}^{l}u_{h}]\!]\right\|_{\partial K_{i}}}\right),$ (2.12a)
@@
Meanwhile, the DG formulation (2.9) with $\sigma_{i}=\sigma^{entropy}_{i}$, while entropy stable, may still generate spurious oscillations in the presence of strong shocks. To remedy such a drawback, we further incorporate $\sigma_{i}^{jump}$ to suppress oscillations and define $\sigma_{i}$ as given in (2.11). $\sigma_{i}^{jump}$, originally proposed by the OFDG method [32] and slightly modified in [40], measures discontinuity intensity and is defined with
$\left\|[\![\partial_{x}^{l}u_{h}]\!]\right\|_{\partial K_{i}}=\left|[\![\frac{\partial^{l}u_{h}}{\partial x^{l}}]\!]_{x_{i-1/2}}\right|+\left|[\![\frac{\partial^{l}u_{h}}{\partial x^{l}}]\!]_{x_{i+1/2}}\right|.$
The coefficient $c_{f}$ is defined as $c_{f}=c_{0}\max\limits_{x\in K_{i}}\{\left|f^{\prime}(u_{h}(x))\right|\}$, as introduced in [33], ensuring a scale-invariant property. Here, $c_{0}$ is a constant. However, since $\sigma_{i}^{entropy}$ in (2.12b) is not scale-invariant, the proposed technique using (2.11) does not possess such a property.
@@
Here, $A\lesssim B$ means that there exists a constant $c>0$ independent of $h$ such that $A\leq cB$. However, we cannot directly estimate the ratio $F_{i}/E_{i}$ unless there is a strictly positive lower bound of $E_{i}$, which is not the case in practice. Therefore, in the simulation, we set
@@ 2.1.3 Theoretical properties
$\max\left\{\frac{F_{i}}{E_{i}},0\right\}<C\,\sigma_{i}^{jump},\quad\forall K_{i}\in\mathcal{K},$ (2.17)
@@
$\mathrm{RHS}$ $=-\sum_{i}{\left(\hat{f}_{i+1/2}^{D}v_{h}|_{i+1/2}^{-}-\hat{f}_{i+1/2}^{D}v_{h}|_{i+1/2}^{+}\right)}$
@@
$=-\sum_{i}{\frac{1}{2}\alpha_{i+1/2}\left(u_{h}|_{i+1/2}^{+}-u_{h}|_{i+1/2}^{-}\right)\left(v\left(u_{h}|_{i+1/2}^{+}\right)-v\left(u_{h}|_{i+1/2}^{-}\right)\right)}$ (2.20)
@@
$=$ $\sum_{i}\left(\sum_{\alpha=1}^{k+1}\hat{\omega}_{i}^{\alpha}\frac{\partial u_{h}}{\partial t}(\hat{x}^{\alpha}_{i})v_{h}(\hat{x}^{\alpha}_{i})+\mathcal{O}(h^{2k+1})\right)$
$=$ $\sum_{i}\left(\sum_{\alpha=1}^{k+1}\hat{\omega}_{i}^{\alpha}\frac{\partial U(u_{h})}{\partial t}(\hat{x}^{\alpha}_{i})+\mathcal{O}(h^{2k+1})\right)$
@@
Here, $\hat{\omega}_{i}^{\alpha}$ and $\hat{x}^{\alpha}_{i}$ are the Gauss-Lobatto weights and quadrature points in each cell $K_{i}$, with $\hat{x}_{i}^{1}=x_{i-1/2}$ and $\hat{x}^{k+1}_{i}=x_{i+1/2}$. This indicates that the entropy can be considered stable, with only a small, high order growth. Also see [15].
@@
For $\sigma_{i}$, since $u$ is smooth, we have $\left|[\![\partial_{x}^{l}u_{h}]\!]\right|=\left|[\![\partial_{x}^{l}(u-u_{h})]\!]\right|$. According to (2.15), the viscosity coefficient has an upper bound $\sigma_{i}\leq C\sigma_{i}^{jump}$. Hence,
$\sum_{i}{\left(\sigma_{i}\right)^{2}}h^{-1}$ $\lesssim h\sum_{i}{\sum_{l=0}^{k}{\left\|[\![\partial_{x}^{l}u_{h}]\!]\right\|_{\partial K_{i}}^{2}h^{2l}}}$
$=h\sum_{i}{\sum_{l=0}^{k}{\left\|[\![\partial_{x}^{l}\left(u-u_{h}\right)]\!]\right\|_{\partial K_{i}}^{2}h^{2l}}}$
@@
$\sigma_{i,m}^{jump}$ $=c_{f}\left(h\left\|[\![u_{h,m}]\!]\right\|_{\partial K_{i}}+\sum_{l=1}^{k}{l\left(l+1\right)h^{l+1}\left\|[\![\partial_{x}^{l}u_{h,m}]\!]\right\|_{\partial K_{i}}}\right),$
@@
Here, similar to the scalar case, $\sigma_{i}^{jump}$ is defined to measure discontinuity on each component without characteristic decomposition. For the Euler equations, we define the coefficient $c_{f}$ as $c_{f}=c_{0}\max\limits_{x\in K_{i}}\{1/H\}$ as in [40], where $H=(\mathcal{E}+p)/\rho$ denotes the enthalpy, and $c_{0}$ is a constant. Note that $c_{f}$ can also be chosen as the scale-invariant version under the OEDG framework [33]. Again, as $\sigma_{i}^{entropy}$ is not scale-invariant, our scheme does not possess this property. The numerical entropy flux $\hat{F}$ is given as the central flux (2.13) as well. The numerical flux $\hat{\mathbf{f}}=\hat{\mathbf{f}}(\mathbf{u}_{L},\mathbf{u}_{R})$ is split into the central part $\hat{\mathbf{f}}^{C}$ and the diffusive part $\hat{\mathbf{f}}^{D}$. For example, for the Lax-Friedrichs flux,
@@
Following the same approach as in proving Theorem 2.2, we can similarly establish that the scheme (2.31) is entropy stable for the hyperbolic system (2.28) with periodic or compact boundaries.
Theorem 2.4.
If
$\sigma_{i}^{entropy}<C\max\limits_{m}\{\sigma_{i,m}^{jump}\},\quad\forall K_{i}\in\mathcal{K},$ (2.32)
then the NOES-DG scheme (2.31) is entropy stable in the sense of
$\int_{\Omega}{\frac{\partial\mathbf{u}_{h}}{\partial t}\cdot\mathbf{v}_{h}\mathrm{d}x}\leq 0.$ (2.33)
Proof.
For Lax-Friedrichs flux, HLL flux, or HLLC flux, we can write
$\hat{\mathbf{f}}^{D}_{i+1/2}=-S_{i+1/2}(\mathbf{u}_{h}|_{i+1/2}^{+}-\mathbf{u}_{h}|_{i+1/2}^{-}),\quad S_{i+1/2}\geq 0.$
Hence, taking $\mathbf{w}=\mathbf{v}_{h}$ in (2.31) and summing of $i$, we have
$\int_{\Omega}{\frac{\partial\mathbf{u}_{h}}{\partial t}\cdot\mathbf{v}_{h}\mathrm{d}x}\leq$ $-\sum_{i}{\left(\hat{\mathbf{f}}_{i+1/2}^{D}\cdot\mathbf{v}_{h}|_{i+1/2}^{-}-\hat{\mathbf{f}}_{i-1/2}^{D}\cdot\mathbf{v}_{h}|_{i-1/2}^{+}\right)}$
$=$ $-\sum_{i}{S_{i+1/2}\left(\mathbf{u}_{h}|_{i+1/2}^{+}-\mathbf{u}_{h}|_{i+1/2}^{-}\right)\cdot\left(\mathbf{v}_{h}|_{i+1/2}^{+}-\mathbf{v}_{h}|_{i+1/2}^{-}\right)}$
$=$ $-\sum_{i}{S_{i+1/2}\left(\mathbf{u}_{h}|_{i+1/2}^{+}-\mathbf{u}_{h}|_{i+1/2}^{-}\right)\cdot\left(\mathbf{v}\left(\mathbf{u}_{h}|_{i+1/2}^{+}\right)-\mathbf{v}\left(\mathbf{u}_{h}|_{i+1/2}^{-}\right)\right)}$
$=$ $-\sum_{i}{S_{i+1/2}\int_{0}^{1}{\left(\mathbf{u}_{h}|_{i+1/2}^{+}-\mathbf{u}_{h}|_{i+1/2}^{-}\right)^{T}}}\mathbf{A}\left(\mathbf{u}_{\theta}\right)\left(\mathbf{u}_{h}|_{i+1/2}^{+}-\mathbf{u}_{h}|_{i+1/2}^{-}\right)\mathrm{d}\theta$
$\leq$ $\ 0$
Here, $\mathbf{u}_{\theta}=\mathbf{u}_{h}|_{i+1/2}^{-}+\theta(\mathbf{u}_{h}|_{i+1/2}^{+}-\mathbf{u}_{h}|_{i+1/2}^{-})$.
@@ 2.3 Two-dimensional system
$\sigma_{K,m}^{jump}$ $=c_{f}\left(h_{K}\left\|[\![u_{h,m}]\!]\right\|_{\partial K}+\sum_{\left|l\right|=1}^{k}{\left|l\right|\left(\left|l\right|+1\right)h_{K}^{\left|l\right|+1}\left\|[\![\partial^{l}u_{h,m}]\!]\right\|_{\partial K}}\right),\ l\in\mathbb{N}^{d},$